A small farm has sheep and chickens. There are twice as
many chicken as sheep, and there are 104 legs between the sheep and the chickens. How many chickens are there?

Answer by stanbon(75887)

A small farm has sheep and chickens. There are twice as
many chicken as sheep, and there are 104 legs between the sheep and the chickens. How many chickens are there?
Equations:
c = 2s
2c+4s = 104 legs
----------------------
Substitute for "c" and solve for "s":
2(2s) + 4s = 104
8s = 104
s = 13 (# of sheep)
----
Solve for "c":
c = 2s = 2*13 = 26 (# of chickens)
----------------
Cheers,
Stan H.
-------------